water flows through a pipe with a cross-sectional area of ​​2 dm ^ 2 which forms a 90 degree bend. the speed of the water in and out of the pipe section is 5cm / s (over the entire cross-sectional area). specify the size and direction of the external force needed to hold the pipe still?


1
Expert's answer
2021-04-26T17:01:49-0400
"F=0.5\sqrt{2}\rho Av^2\\F=0.5\sqrt{2}(1000)(0.02)(0.05)^2=0.035\ N\\\theta=45\degree"